OpenStudy (anonymous):

I am so brain dead on how to integrate sec(2x)dx. I converted it to the integration of 1/(cos^2(x)-sin^2(x)) and I am so blank on how to go on from here.

OpenStudy (anonymous):

\[\int\sec 2x~dx\] Let \(u=2x\), so \(\dfrac{1}{2}du=dx\). \[\frac{1}{2}\int \sec u~du\] The next step isn't incredibly intuitive, so it helps to recall your trigonometric derivatives: \[\frac{1}{2}\int \sec u\cdot\frac{\sec u+\tan u}{\sec u+\tan u}~du\] When you distribute the \(\sec u\), your numerator will be \(\sec^2u+\sec u\tan u\), which is the derivative of the denominator, and so you can use another substitution to proceed.